OpenStudy (anonymous):

- tan^2x + sec^2x = 1 Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

OpenStudy (zehanz):

Identities are: \(\tan x=\dfrac{\sin x}{\cos x}\) and \(\sec x = \dfrac{1}{\cos x}\)

OpenStudy (zehanz):

This "identity" doesn't seem to be true, however... Are you sure this is the way it was stated?

OpenStudy (anonymous):

\[-\tan^2x + \sec^2x=1\]

OpenStudy (zehanz):

Sorry, I misread it, now it is much clearer!

OpenStudy (anonymous):

so wouldn't i just move -tan^2x to the right side ?

OpenStudy (zehanz):

So after using the identities, I get: \(\dfrac{-\sin^2x}{\cos^2x}+\dfrac{1}{\cos^2x}\) Write it as one fraction, then use sin²x+cos²x=1

OpenStudy (anonymous):

isnt this a trig proof?

OpenStudy (zehanz):

I wouldn't move -tan²x to the right side, because you have to show that the left side is equal to 1. I would just rework the left side until it's obviously equal to 1.

OpenStudy (zehanz):

I mean this: the left side can now be written as: \(\dfrac{1-\sin^ 2x}{\cos^2x}\). If you use that sin²x+cos²x=1, you can change the numerator to something much simpler...

OpenStudy (anonymous):

what do i do after that

OpenStudy (zehanz):

Well, because sin²x+cos²x=1, cos²x=1-sin²x, so you can replace the numerator with cos²x.

OpenStudy (zehanz):

So you end up with \(\dfrac{\cos^2x}{\cos^2x}\)=...

OpenStudy (anonymous):

thank you :)

OpenStudy (zehanz):

yw!